# One Subgoal at a Time: Zero-Shot Generalization to Arbitrary Linear Temporal Logic Requirements in Multi-Task Reinforcement Learning

### v1

@@ Acknowledgement
This work was supported in part by the U.S. National Science Foundation under grant CCF-2340776.
@@ A.1 Environments
[…]
We exclude the FlatWorld[23, 44] environment from our evaluation. It contains colored regions similar to ZoneEnv. However, the region layout is fixed, meaning that certain Boolean formulas over the atomic propositions cannot be evaluated since they are never true in this environment. Additionally, the observation space is limited to the agent’s $(x,y)-$position. Due to this lack of variability and expressiveness, we exclude FlatWorld from our evaluation.
@@ A.2 LTL Specifications
We evaluate zero-shot generalization across a range of specifications, including complex and nested finite-horizon tasks (Table 3), complex infinite-horizon tasks (Table 4), varying specification complexity with different sequence lengths of reach-only and reach-avoid specifications (Table 5), varying environment complexity in reach-avoid tasks (Table 5), and different numbers of atopic propositions (Table 6). The specifications $\varphi_{1}-\varphi_{4}$ and $\varphi_{9}-\varphi_{12}$ are from [23]. To better evaluate each method’s generalization ability and its handling of the trade-off between task progression and safety constraints, we construct more complex specifications, $\varphi_{5}$$\varphi_{8}$ and $\varphi_{13}$$\varphi_{16}$, which feature longer temporal sequences and stricter safety requirements, i.e., more atomic propositions appear under negation.
Table 3: Complex finite-horizon specifications used in our evaluation. Results are shown in Table 1.
LetterWorld
$\varphi_{1}$ $\mathsf{F}\,\left(\mathsf{a}\land(\neg\mathsf{b}\ \;\mathsf{U}\;\ \mathsf{c})\right)\land\mathsf{F}\,\mathsf{d}$
$\varphi_{2}$ $(\mathsf{F}\,\mathsf{d})\land\left(\neg\mathsf{f}\ \;\mathsf{U}\;\ (\mathsf{d}\land\mathsf{F}\,\mathsf{b})\right)$
$\varphi_{3}$$\neg\mathsf{a}\ \;\mathsf{U}\;\ \left(\mathsf{b}\land\left(\neg\mathsf{c}\ \;\mathsf{U}\;\ \left(\mathsf{d}\land(\neg\mathsf{e}\ \;\mathsf{U}\;\ \mathsf{f})\right)\right)\right)$
$\varphi_{4}$ $\left(\mathsf{a}\lor\mathsf{b}\lor\mathsf{c}\lor\mathsf{d}\Rightarrow\mathsf{F}\,\left(\mathsf{e}\land\mathsf{F}\,\left(\mathsf{f}\land\mathsf{F}\,\mathsf{g}\right)\right)\right)\ \;\mathsf{U}\;\ \left(\mathsf{h}\land\mathsf{F}\,\mathsf{i}\right)$
$\varphi_{5}$ $\mathsf{F}\,\left(\mathsf{d}\land\left(\neg(\mathsf{a}\lor\mathsf{b})\;\mathsf{U}\;(\mathsf{b}\land(\neg\mathsf{e}\;\mathsf{U}\;\mathsf{c}))\right)\right)\land\mathsf{F}\,\left(\neg(\mathsf{f}\lor\mathsf{g}\lor\mathsf{h})\;\mathsf{U}\;\mathsf{a}\right)$
$\varphi_{6}$ $\mathsf{F}\,\left(\left(\mathsf{k}\land\left(\left(\neg\mathsf{b}\lor\mathsf{c}\right)\;\mathsf{U}\;\mathsf{f}\right)\right)\land\left(\neg(\mathsf{a}\lor\mathsf{e}\lor\mathsf{h})\;\mathsf{U}\;\mathsf{g}\right)\right)\land\mathsf{F}\,\mathsf{d}$
$\varphi_{7}$$\neg(\mathsf{j}\lor\mathsf{b}\lor\mathsf{d})\ \;\mathsf{U}\;\ \left(\mathsf{a}\land(\neg\mathsf{c}\;\mathsf{U}\;(\mathsf{f}\land\mathsf{F}\,(\mathsf{g}\land(\neg\mathsf{d}\;\mathsf{U}\;\mathsf{e}))))\right)$
$\varphi_{8}$$\neg(\mathsf{f}\lor\mathsf{g})\ \;\mathsf{U}\;\ \left(\mathsf{a}\land\left(\neg\mathsf{b}\;\mathsf{U}\;\mathsf{c}\right)\land\mathsf{F}\,\left(\mathsf{d}\land\left(\neg\mathsf{e}\;\mathsf{U}\;\mathsf{f}\right)\right)\right)$
ZoneEnv
$\varphi_{9}$ $(\mathsf{F}\,\mathsf{blue})\land\left(\neg\mathsf{blue}\ \;\mathsf{U}\;\ (\mathsf{green}\land\mathsf{F}\,\mathsf{yellow})\right)$
$\varphi_{10}$$\neg(\mathsf{magenta}\lor\mathsf{yellow})\ \;\mathsf{U}\;\ \left(\mathsf{blue}\land\mathsf{F}\,\mathsf{green}\right)$
$\varphi_{11}$$\neg\mathsf{green}\ \;\mathsf{U}\;\ \left((\mathsf{blue}\lor\mathsf{magenta})\land(\neg\mathsf{green}\ \;\mathsf{U}\;\ \mathsf{yellow})\right)$
$\varphi_{12}$ $\left(\mathsf{green}\lor\mathsf{blue}\Rightarrow\left(\neg\mathsf{yellow}\ \;\mathsf{U}\;\ \mathsf{magenta}\right)\right)\ \;\mathsf{U}\;\ \mathsf{yellow}$
$\varphi_{13}$ $\mathsf{F}\,\left(\mathsf{green}\land\left(\neg(\mathsf{blue}\lor\mathsf{yellow})\;\mathsf{U}\;(\mathsf{yellow}\land(\neg\mathsf{magenta}\;\mathsf{U}\;\mathsf{blue}))\right)\right)\land\mathsf{F}\,\left(\neg\mathsf{green}\;\mathsf{U}\;\mathsf{yellow}\right)$
$\varphi_{14}$ $\mathsf{F}\,\left((\mathsf{blue}\lor\mathsf{green})\land\left(\neg\mathsf{yellow}\;\mathsf{U}\;(\mathsf{blue}\land(\neg(\mathsf{green}\lor\mathsf{magenta})\;\mathsf{U}\;\mathsf{magenta}))\right)\right)\land\mathsf{F}\,\left(\mathsf{yellow}\land(\neg\mathsf{blue}\;\mathsf{U}\;\mathsf{green})\right)$
$\varphi_{15}$$\neg(\mathsf{magenta}\lor\mathsf{yellow})\ \;\mathsf{U}\;\ \left(\mathsf{blue}\land(\neg\mathsf{green}\;\mathsf{U}\;(\mathsf{yellow}\land\mathsf{F}\,(\mathsf{green}\land(\neg\mathsf{blue}\;\mathsf{U}\;\mathsf{magenta}))))\right)$
$\varphi_{16}$ $\mathsf{F}\,\left(\mathsf{blue}\land\left(\neg\mathsf{yellow}\;\mathsf{U}\;(\mathsf{green}\land\mathsf{F}\,(\mathsf{yellow}\land(\neg(\mathsf{magenta}\lor\mathsf{green})\;\mathsf{U}\;\mathsf{blue})))\right)\right)$
Table 4: Infinite-horizon specifications used in our evaluation. Results are shown in Table 2.
LetterWorld $\psi_{1}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{blue}\land\mathsf{G}\,\mathsf{F}\,\mathsf{green}\land\mathsf{G}\,\neg(\mathsf{yellow}\lor\mathsf{magenta})$
$\psi_{2}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{blue}\land\mathsf{G}\,\mathsf{F}\,\mathsf{yellow}\land\mathsf{G}\,\mathsf{F}\,\mathsf{green}\land\mathsf{G}\,\neg\mathsf{magenta}$
$\psi_{3}$ $\mathsf{F}\,\mathsf{G}\,\mathsf{yellow}\land\mathsf{G}\,\neg(\mathsf{green}\lor\mathsf{blue}\lor\mathsf{magenta})$
ZoneEnv $\psi_{4}$ $\mathsf{G}\,\mathsf{F}\,(\mathsf{e}\land(\neg\mathsf{a}\ \;\mathsf{U}\;\ \mathsf{f}))\land\mathsf{G}\,\neg(\mathsf{c}\lor\mathsf{d})$
$\psi_{5}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{a}\land\mathsf{G}\,\mathsf{F}\,\mathsf{b}\land\mathsf{G}\,\mathsf{F}\,\mathsf{c}\land\mathsf{G}\,\neg(\mathsf{e}\lor\mathsf{f}\lor\mathsf{i})$
$\psi_{6}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{c}\land\mathsf{G}\,\mathsf{F}\,\mathsf{a}\land\mathsf{G}\,\mathsf{F}\,(\mathsf{e}\land(\neg\mathsf{f}\ \;\mathsf{U}\;\ \mathsf{g}))\land\mathsf{G}\,\mathsf{F}\,\mathsf{k}\land\mathsf{G}\,\neg(\mathsf{i}\lor\mathsf{j})$
@@
n=2 ($(\neg\mathsf{blue}\land\neg\mathsf{green})\ \;\mathsf{U}\;\ \mathsf{yellow}$
($(\neg\mathsf{yellow}\land\neg\mathsf{magenta})\ \;\mathsf{U}\;\ \mathsf{green}$
($(\neg\mathsf{green}\land\neg\mathsf{magenta})\ \;\mathsf{U}\;\ \mathsf{blue}$
@@ Appendix B Implementation Details
We use the official codebases of the baselines: LTL2Action [^1] https://github.com/LTL2Action/LTL2Action, GCRL-LTL [^2] https://github.com/RU-Automated-Reasoning-Group/GCRL-LTL, and DeepLTL [^3] https://github.com/mathiasj33/deep-ltl. Our implementation is adapted from DeepLTL by incorporating a safety value function and a state-dependent Lagrangian multiplier and removing DeepLTL-specific modules. For DeepLTL, LTL2Action, and our method, we use a fully connected actor network with three hidden layers of sizes [64, 64, 64]. The value function network has two hidden layers of sizes [64, 64]. For our method, both the safety value function and the Lagrangian multiplier network share the same architecture as the value function. GCRL-LTL employs a larger actor-critic architecture with [512, 1024, 256] units due to its use of a graph neural network. For fair comparison, we use the Adam optimizer with a learning rate of $3e-4$ and train all methods for 15M environment interactions. The discount factor is set to $\gamma=0.94$ for LetterWorld and $\gamma=0.998$ for ZoneEnv. Additional hyperparameter details are provided in the code in the supplementary materials. We train all the methods using a 16-core AMD CPU and an NVIDIA GeForce RTX 4090 GPU, and it takes roughly 3 hours to train our model.
Our method is summarized in Algorithm 1. We begin by constructing a subgoal set that includes all feasible subgoals: we first enumerate each assignment $\alpha\in 2^{AP}$ as a candidate $\alpha^{+}$. For each such $\alpha^{+}$, we filter out the remaining assignments that conflict with it. We then enumerate all possible combinations of the filtered assignments to form $A^{-}$, resulting in the full set of feasible reach-avoid subgoals $\xi=\{(\alpha^{+},A^{-})_{i}\}_{i=1}^{M}$, from which subgoals are sampled uniformly. The agent receives a reward $r=1$ if the subgoal is satisfied and $r=0$ otherwise. At each timestep, it receives a safety signal $h=-1$ if the subgoal is not violated, and $h=1$ otherwise. The reward and safety functions are defined as $r=\mathbbm{1}[L(s^{\sigma})\in\alpha^{+}]$ and $h=2\cdot\mathbbm{1}[L(s^{\sigma})\in A^{-}]-1$, respectively. If the current subgoal is satisfied, we sample a new subgoal whose avoid set does not contain the current state, i.e., $L(s_{t})\notin A^{-}$, and whose reach set also excludes the current state, i.e., $L(s_{t})\notin\alpha^{+}$, to prevent the new subgoal from being immediately violated or trivially satisfied. If a subgoal is violated, the episode terminates. We use Generalized Advantage Estimation (GAE) [41] to estimate value functions:
@@
1:Initial policy $\pi$, value and safety value function $V_{r}$, $V_{h}$, state-dependent Lagrangian multiplier $\lambda$, subgoal set $\xi$, maximum training iterations $K$, interactions per iteration $N$
@@
[…]
@@
8:   $\mathcal{Q}_{o}\leftarrow\mathcal{Q}_{c}$triangleright Remember the old states.
9:   $\mathcal{Q}_{c}\leftarrow\{q^{\prime}\mid\exists q\in\mathcal{Q}_{o},(q,\alpha,q^{\prime})\in\delta\}$triangleright Update the current set of states.
@@
13:     $\mathcal{U}\leftarrow\mathcal{U}\cup\{(q,\alpha^{+})\mid q\in\left(\mathcal{Q}_{c}-\mathcal{F}\right)\}$triangleright If $q$ is not accepting, mark this subgoal as unsatisfiable.
@@ Appendix C Additional Experiments
$\eta_{v}\downarrow$ $0.02_{\pm 0.01}$ $0.02_{\pm 0.01}$ $0.03_{\pm 0.01}$ $0.02_{\pm 0.01}$
$\eta_{o}\downarrow$ $0.02_{\pm 0.02}$ $0.04_{\pm 0.02}$ $0.06_{\pm 0.03}$ $0.05_{\pm 0.02}$
$\mu\downarrow$ $306.58_{\pm 16.88}$ $314.88_{\pm 17.02}$ $313.10_{\pm 10.61}$ $323.30_{\pm 12.57}$
[…]
@@
Figure 14: Illustration of ZoneEnv under increasing environment complexity, including different zone sizes and different numbers of zones per atomic proposition. The default zone size is 0.4, and the default number of zones per atomic proposition is 2. The trajectories are generated by our method and correspond to the evaluated specification ($(\neg\mathsf{blue}\land\neg\mathsf{green})\ \;\mathsf{U}\;\ \mathsf{yellow}$.

### v6

@@ Acknowledgement
The authors thank the anonymous reviewers for their invaluable feedback and constructive suggestions. This work was supported in part by the U.S. National Science Foundation under grant CCF-2340776.
Broader Impacts
Our work aims to improve the generalization of reinforcement learning to satisfy arbitrary LTL specifications. This framework enables flexible and expressive task definitions, but it also places responsibility on the user to ensure that the specifications are well designed. The behavior of the agent is tightly coupled with the given specification, so poorly constructed or harmful specifications may lead to unsafe or unintended behaviors. Thus, careful design and verification of specifications are essential when applying such methods in real-world settings.
@@ A.1 Environments
We use the LetterWorld [vaezipoor2021ltl2action] and ZoneEnv [vaezipoor2021ltl2action, jackermeierdeepltl] for evaluation. The LetterWorld is a $7\times 7$ grid world that contains 12 letters corresponding to atomic propositions $AP=\{\mathsf{a},\mathsf{b},\dots,\mathsf{l}\}$. Each letter appears twice and is randomly placed on the grid. The observation consists of the entire grid in an egocentric view. The agent can move in four directions: up, down, left, and right. The grid wraps around, i.e., if the agent moves out of bounds, it reappears on the opposite side. The maximum episode length is $T=75$. The ZoneEnv contains different colored regions corresponding to atomic propositions $AP=\{\mathsf{blue},\mathsf{green},\mathsf{magenta},\mathsf{yellow}\}$ and walls acting as boundaries. Our implementation is adapted from [jackermeierdeepltl]. We use the point robot that has a continuous action space for acceleration and steering. Observations consist of Lidar measurements of the colored regions and ego-centric state information from other sensors. The maximum episode length is $T=1000$. We further construct several variants of the ZoneEnv to evaluate the generalization performance of our method and baselines by modifying: (1) the region size, with $r=\{0.2,0.3,0.4(\text{default}),0.5,0.6\}$; (2) the number of regions per atomic proposition, with $\text{count}=\{1,2(\text{default}),3,4,5\}$; (3) and the number of atomic propositions, expanded to include new symbols $\{\mathsf{red},\mathsf{cyan},\mathsf{orange},\mathsf{purple},\mathsf{teal},\mathsf{lime}\}$. Illustrations of the environments are provided in Figure 8 and trajectories under different LTL specifications are provided in Figure 12, 15, and 16.
We exclude the FlatWorld[jackermeierdeepltl, shah2024LTLConstrained] environment from our evaluation. It contains colored regions similar to ZoneEnv. However, the region layout is fixed, meaning that certain Boolean formulas over the atomic propositions cannot be evaluated since they are never true in this environment. Additionally, the observation space is limited to the agent’s $(x,y)$-position. Due to this lack of variability and expressiveness, we exclude FlatWorld from our evaluation.
@@ A.2 LTL Specifications
We evaluate zero-shot generalization across a range of specifications, including complex and nested finite-horizon tasks (Table 1), complex infinite-horizon tasks (Table 1), varying specification complexity with different sequence lengths of reach-only and reach-avoid specifications (Table 4), varying environment complexity in reach-avoid tasks (Table 4), and different numbers of atopic propositions (Table 5). The specifications $\varphi_{1}-\varphi_{4}$ and $\varphi_{9}-\varphi_{12}$ are from [jackermeierdeepltl]. To better evaluate each method’s generalization ability and its handling of the trade-off between task progression and safety constraints, we construct more complex specifications, $\varphi_{5}$$\varphi_{8}$ and $\varphi_{13}$$\varphi_{16}$, which feature longer temporal sequences and stricter safety requirements, i.e., more atomic propositions appear under negation.
@@
n=2 $(\neg\mathsf{blue}\land\neg\mathsf{green})\ \;\mathsf{U}\;\ \mathsf{yellow}$
$(\neg\mathsf{yellow}\land\neg\mathsf{magenta})\ \;\mathsf{U}\;\ \mathsf{green}$
$(\neg\mathsf{green}\land\neg\mathsf{magenta})\ \;\mathsf{U}\;\ \mathsf{blue}$
@@ Appendix B Implementation Details
We use the official codebases of the baselines: LTL2Action [^1] https://github.com/LTL2Action/LTL2Action, GCRL-LTL [^2] https://github.com/RU-Automated-Reasoning-Group/GCRL-LTL, RAD-embeddings [^3] https://github.com/RAD-Embeddings/neurips24, and DeepLTL [^4] https://github.com/mathiasj33/deep-ltl. Our implementation is adapted from DeepLTL by incorporating a safety value function and a state-dependent Lagrangian multiplier and removing DeepLTL-specific modules. For DeepLTL, LTL2Action, and our method, we use a fully connected actor network with three hidden layers of sizes [64, 64, 64]. The value function network has two hidden layers of sizes [64, 64]. For our method, both the safety value function and the Lagrangian multiplier network share the same architecture as the value function. GCRL-LTL employs a larger actor-critic architecture with [512, 1024, 256] units due to its use of a graph neural network. For fair comparison, we use the Adam optimizer with a learning rate of 3e43e-4 and train all methods for 15M environment interactions. The discount factor is set to $\gamma=0.94$ for LetterWorld and $\gamma=0.998$ for ZoneEnv. Additional hyperparameter details are provided in the code in the supplementary materials. We train all the methods using a 16-core AMD CPU and an NVIDIA GeForce RTX 4090 GPU, and it takes roughly 3 hours to train our model.
Our method is summarized in Algorithm 1. We begin by constructing a subgoal set that includes all feasible subgoals: we first enumerate each assignment $\alpha\in 2^{AP}$ as a candidate $\alpha^{+}$. For each such $\alpha^{+}$, we filter out the remaining assignments that conflict with it. We then enumerate all possible combinations of the filtered assignments to form $A^{-}$, resulting in the full set of feasible reach-avoid subgoals $\xi=\{(\alpha^{+},A^{-})_{i}\}_{i=1}^{M}$, from which subgoals are sampled uniformly. The agent receives a reward $r=1$ if the subgoal is satisfied and $r=0$ otherwise. At each timestep, it receives a safety signal h=1h=-1 if the subgoal is not violated, and $h=1$ otherwise. The reward and safety functions are defined as $r=\mathbbm{1}[L(s^{\sigma})\in\alpha^{+}]$ and $h=2\cdot\mathbbm{1}[L(s^{\sigma})\in A^{-}]-1$, respectively. If the current subgoal is satisfied, we sample a new subgoal whose avoid set does not contain the current state, i.e., $L(s_{t})\notin A^{-}$, and whose reach set also excludes the current state, i.e., $L(s_{t})\notin\alpha^{+}$, to prevent the new subgoal from being immediately violated or trivially satisfied. If a subgoal is violated, the episode terminates. We use Generalized Advantage Estimation (GAE) [schulman2015high] to estimate value functions:
@@
1:Initial policy $\pi$, value and safety value function $V_{r}$, $V_{h}$, state-dependent Lagrangian multiplier $\lambda$, subgoal set $\xi$, maximum training iterations $K$, interactions per iteration NN
@@
For Lidar measurements in the ZoneEnv, $f(\cdot)$ is implemented as the element-wise $\min$ to obtain the closest distance along each Lidar beam, such that the resulting $s^{\sigma}$ For grid maps in LetterWorld, we assign distinct values to each cell associated with $\sigma$ at location $(i,j)$: $f(s_{i,j})=v_{\text{reach}}$ if $L(s_{i,j})=\alpha^{+}$; $v_{\text{avoid}}$ if $L(s_{i,j})\in A^{-}$; $v_{\text{neutral}}$ otherwise. This reduction allows the agent to focus on "reach" and "avoid" rather than specific set of atomic propositions and eliminates the need of encoding subgoals, enabling efficient policy learning and generalization to new atomic propositions that satisfies the conditioned discussed in Section 4.2 as shown later in Appendix C. If multiple observation functions are available, each capable of observing an assignment $\alpha\in 2^{AP}$, we can first fuse the observations within each observation type based on the current subgoal $\sigma$ and then concatenate the results to form the final state $s^{\sigma}$. In cases where each atomic proposition is tied to a distinct observation type, our method can be applied; however, learning policies in such settings is inherently intractable due to the exponential number of subgoal-state combinations.
@@
8:   $\mathcal{Q}_{o}\leftarrow\mathcal{Q}_{c}$ $\triangleright$ Remember the old states.
9:   $\mathcal{Q}_{c}\leftarrow\{q^{\prime}\mid\exists q\in\mathcal{Q}_{o},(q,\alpha,q^{\prime})\in\delta\}$ $\triangleright$ Update the current set of states.
@@
13:     $\mathcal{U}\leftarrow\mathcal{U}\cup\{(q,\alpha^{+})\mid q\in\left(\mathcal{Q}_{c}-\mathcal{F}\right)\}$ $\triangleright$ If $q$ is not accepting, mark this subgoal as unsatisfiable.
@@ Appendix C Additional Experiments
$\eta_{v}\downarrow$ $02_{\pm 0.01} 0.0$$2_{\pm 0.01} 0.0$$3_{\pm 0.01} 0.0$$02_{\pm 0.01}$
$\eta_{o}\downarrow$ $02_{\pm 0.02} 0.0$$4_{\pm 0.02} 0.0$$6_{\pm 0.03} 0.0$$05_{\pm 0.02}$
$\mu\downarrow$ $306.58_{\pm 16.88}$ $4.88_{\pm 17.02} 313.1$$3.10_{\pm 10.61}$ $323.30_{\pm 12.57}$
Figure 12: Illustration of ZoneEnv with an increasing number of atomic propositions. The trajectories shown correspond to the first specification for $|\text{AP}|=4,6,8,10$ in Table 5.
@@
Table 7: Performance in ZoneEnv with Ant agent. We report the success rate $\eta_{s}$, violation rate $\eta_{v}$, and the average steps to satisfy the specifications $\mu$. Each value is averaged over 5 seeds, with 100 trajectories per seed.
GenZ-LTL(ours) DeepLTL
$\eta_{s}\uparrow$ $\eta_{v}\downarrow$ $\mu\downarrow$ $\eta_{s}\uparrow$ $\eta_{v}\downarrow$ $\mu\downarrow$
ZoneEnv
$\varphi_{9}$ $97_{\pm 0.02} 0.0$$02_{\pm 0.01}$ $305.32_{\pm 46.50}$ $0.00_{\pm 0.01}$ $0.23_{\pm 0.08}$ $865.50_{\pm 72.83}$
$\varphi_{10}$ $91_{\pm 0.03} 0.0$$09_{\pm 0.03}$ $202.00_{\pm 53.33}$ $0.03_{\pm 0.02}$ $0.39_{\pm 0.11}$ $588.58_{\pm 166.33}$
$\varphi_{11}$ $95_{\pm 0.02} 0.0$$04_{\pm 0.02}$ $182.88_{\pm 38.66}$ $0.07_{\pm 0.02}$ $0.27_{\pm 0.05}$ $604.87_{\pm 28.22}$
$\varphi_{12}$ $98_{\pm 0.01} 0.0$$00_{\pm 0.00}$ $125.97_{\pm 35.66}$ $24_{\pm 0.05} 0.0$$05_{\pm 0.04}$ $392.04_{\pm 64.39}$
$\varphi_{9-12}$ $95_{\pm 0.04} 0.0$$04_{\pm 0.04}$ $204.04_{\pm 77.69}$ $0.09_{\pm 0.10}$ $0.23_{\pm 0.14}$ $568.15_{\pm 172.57}$
$\varphi_{13}$ $93_{\pm 0.04} 0.0$$00_{\pm 0.00}$ $459.99_{\pm 86.69}$ $_{\pm 0.00} 0.00±0.00$$_{\pm 0.00}$ -
$\varphi_{14}$ $95_{\pm 0.02} 0.0$$00_{\pm 0.00}$ $434.41_{\pm 76.87}$ $_{\pm 0.00} 0.00±0.00$$_{\pm 0.00}$ -
$\varphi_{15}$ $89_{\pm 0.02} 0.0$$08_{\pm 0.02}$ $429.76_{\pm 75.75}$ $0.00_{\pm 0.00}$ $0.39_{\pm 0.13}$ -
$\varphi_{16}$ $95_{\pm 0.03} 0.0$$00_{\pm 0.00}$ $449.89_{\pm 83.26}$ $_{\pm 0.00} 0.00±0.00$$_{\pm 0.00}$ -
$\varphi_{13-16}$ $93_{\pm 0.04} 0.0$$02_{\pm 0.04}$ $443.51_{\pm 73.31}$ $0.00_{\pm 0.00}$ $0.10_{\pm 0.18}$ -
Figure 14: Visualization of the Ant agent in ZoneEnv.
Complex Dynamics. To evaluate the performance of our method in environments with more complex dynamics, we conduct additional experiments using the Ant agent [ji2023safety], consisting of a torso and four legs connected by hinge joints, in the Zone environment. The state space includes the agent’s ego-state (40 dimensions, compared to 12 for the Point agent used in our main experiments) and LiDAR observations of the regions. A visualization of the Ant agent is shown in Figure 14. The action space has 8 dimensions, corresponding to the torques applied to the Ant’s joints to coordinate leg movements (compared to 2 dimensions for the Point agent). We train and evaluate our method against DeepLTL, the strongest baseline in our main results. For a fair comparison, all training procedures, training parameters, and neural network sizes were kept the same (note that the last one in particular could result in some performance drop as the state dimension and action dimension are bigger for the ant agent). Both methods are trained until convergence. The evaluation results are shown in Table 7. We can observe that our method consistently outperforms the baseline across all metrics, demonstrating an even larger performance gain in environments with more complex agent dynamics. Note that for the Ant agent, in addition to the LTL specifications, there is a built-in safety constraint that the agent should not fall headfirst, enforced as a hard constraint with immediate episode termination. Although a negative reward is assigned in such cases, as is done when a specification is violated for DeepLTL, it does not explicitly model safety constraints as our method does through HJ reachability, making it less effective in handling such cases. This impedes the learning of coordinated locomotion and then the satisfaction of LTL specifications, leading to degraded performance.
@@
Figure 16: Illustration of ZoneEnv under increasing environment complexity, including different zone sizes and different numbers of zones per atomic proposition. The default zone size is 0.4, and the default number of zones per atomic proposition is 2. The trajectories are generated by our method and correspond to the evaluated specification $(\neg\mathsf{blue}\land\neg\mathsf{green})\ \;\mathsf{U}\;\ \mathsf{yellow}$.